OpenStudy (anonymous):

What is the y-intercept of the graph of y = 4x - 2?

OpenStudy (anonymous):

the y intercept is -2 obviously......since it is of form y=mx+c

OpenStudy (anonymous):

1/16 8 0 1

OpenStudy (anonymous):

can't do -2 so wtf>?

OpenStudy (anonymous):

sry i didnt get u?

OpenStudy (anonymous):

i can not pic -2

OpenStudy (anonymous):

ho....u can point as (0,-2)

OpenStudy (anonymous):

i post ss

OpenStudy (anonymous):

oooooooooooooooh

OpenStudy (anonymous):

that is not \[y=4x-2\] that is \[y=4^x-2\]

OpenStudy (anonymous):

hope u understood my intention

OpenStudy (anonymous):

set equal 0 and solve via \[4^x-2=0\] \[4^x=2\] \[x=\frac{1}{2}\] since \[4^{\frac{1}{2}}=\sqrt{4}=2\]

OpenStudy (anonymous):

answer is 1/2

OpenStudy (anonymous):

dude i can't do 1/2

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

well that is because i found the x- intercept. for the y - intercept replace x by 0 and see what you get

OpenStudy (anonymous):

unfortunately you get \[y=4^0-2=1-2=-1\] and i don't see that answer either!

OpenStudy (anonymous):

So ether we are to stupid to get it right or my school is to stupid to wright something that can be done